OpenStudy (anonymous):

While in a store, your father purchases drinks for 6 people. Part of your family wants coffee and the rest want a soda. Coffee cost $.49 per cup and sodas are $1.29 each. The cost of the drinks before tax was $6.14. (a).Write a mathematical model that represents the total # cost of cups of coffee & sodas. (b). Write a mathematical model to represent the cost of coffee & soda. (c).Solve the system of equations using the elimination method.

OpenStudy (anonymous):

Let c and s be the number of coffee cups and sodas respectively. $ will be expressed as pennies. Solve the following for c and s: c+s=6 49c+129s=614 c=6-s 49(6-s)+129s=614 //Expand 294 + 80 s = 614 80 s = 320 s=4 c=6-4 c=2